OpenStudy (anonymous):

Use logarithmic differentiation to calculate the derivative: f(x) = (1+sin(x))^(1/x)

OpenStudy (anonymous):

I know you take the ln of both sides, but I get stuck because I don't know what to do with the power!! Help? Pretty please?

hartnn (hartnn):

\(\log x^n = n \log x\) \(\log (1+\sin x)^{1/x} = (1/x) \log (1+\sin x)\)

OpenStudy (anonymous):

Okay, I got that far, but I'm stuck!

OpenStudy (anonymous):

...unless you use the product rule, which I think you don't do...

hartnn (hartnn):

you have to use product rule, and ofcourse chain rule for log(1+sin x)

OpenStudy (anonymous):

Okay, just a sec while I work this out then...

hartnn (hartnn):

sure, take your time and tell me what u get ?

OpenStudy (anonymous):

Final answer: \[(1+\sin(x))^{1/x}*(\frac{ \cos(x)}{(x)(1+\sin(x)} -\frac{ \ln(\sin(x) }{ x ^{2} })\] I'm giving the equation doodad down there a whirl, forgive any errors! I got the answer in my book, too, so it's right! Thanks for all the help, we literally just learned this yesterday!

hartnn (hartnn):

welcome ^_^ you learned fast! good :)

OpenStudy (anonymous):

Thanks, I pick things up pretty quick! And I love your avatar/icon!

hartnn (hartnn):

me too :D